# Prompting Techniques for Secure Code Generation: A Systematic Investigation

### v1

@@ 8. Threats to Validity
Construct Validity. As mentioned in Section 5.3, Bandit (as many static analysis tools alike) may output false positive results. To address this threat, we performed a manual validation of Bandit output over a small sample of GPT-3-generated code snippets. Particularly, we inspected 15 of the 150 snippets (around 10%) produced by the baseline prompt template and compared our assessment against the one obtained with Bandit. Overall, although we encountered 3 false negative cases, we could manually confirm all weaknesses flagged by Bandit. These findings suggest that the tool is sufficiently reliable for a comparative analysis albeit with a minor margin of error. The validity analysis of code responses generated by all LLMs was conducted by a single author, potentially introducing biases in the evaluation process. Nonetheless, efforts were made to mitigate such biases by explicitly outlining the criteria for assessing the validity of code snippets, as detailed in Section 5.2. We also acknowledge that the prompting techniques underwent evaluation using prompt templates created by us. These generated templates might have influenced the results obtained for each technique from the LLMs. However, attention was given to crafting the templates, adhering closely to the design and examples that demonstrated optimal results in the respective papers that introduced these techniques.
External Validity. This study evaluates security in Python code which may affect the generalizability of our results to other programming languages such as C/C++ or JavaScript. However, we focused on weaknesses in Python code, given its continued popularity. Moreover, the LLMs used in this study have demonstrated competence in generating functional Python code, which further motivated us to prioritize the evaluation of Python code. Additionally, as already addressed in Section 5.1, this study was conducted only using the OpenAI models. As previously stated, this decision was made due to the popularity of these models in the prompt engineering literature and their demonstrated proficiency in handling coding tasks articulated in natural language, as identified during a preliminary model selection examination by us. It is also worth mentioning that, we focused on prompting techniques that do not rely on demonstrative examples. This choice stemmed from a user study (Perry et al., 2023) which highlighted that users predominantly interacted with AI assistants using natural language coding task specifications or instructions, without supplying demonstrative examples. Furthermore, the LLMSecEval dataset contains NL prompts for only 18 out of Top 25 CWEs of the year 2021. However, 15 out of the 18 CWEs considered in this study have retained their position on the Top 25 list of 2022 and 2023 proving the continued relevance and significance of our research findings.
@@ 9. Conclusion
Our analysis reaffirms the prevalence of security weaknesses in code generated by LLMs when prompted with NL instructions, with significant challenges stemming from CWE-78, CWE-259, CWE-94, and CWE-330. Among the prompting techniques investigated, RCI, a refinement-based approach, exhibited notable effectiveness in preventing security weaknesses in LLM-generated code. Particularly noteworthy was its performance with GPT-4, where it reduced the average weakness density by 77.5% compared to baseline prompting that includes no security specifications. Although RCI demonstrated the highest performance, to the extent of our knowledge, this technique has not been applied for secure code generation in existing literature. This highlights the need for additional research to investigate refinement-based methods, like RCI, which leverage self-critiquing and improvement capabilities of LLMs to enhance security in LLM-generated code. Conversely, persona/memetic proxy techniques, which involve priming the model to adopt a specific role such as that of a software security expert in our case, demonstrated the poorest performance. Zero-shot prompting yielded surprisingly favorable outcomes considering its straightforward nature, performing better than zero-shot CoT and persona/memetic proxy yet falling short of RCI. However, zero-shot prompting holds promise due to its simplicity and relative performance, provided an optimal prompt can be identified.
@@ 10. Replication Package
All data collected and generated in this study are available in https://figshare.com/s/1df4fcedde2901e76870. This repository contains the results of the literature review along with the information on the prompting techniques that were removed from our final selection and the rationale behind this exclusion. Furthermore, the repository contains the code generated by all 3 LLMs for the 7 prompt templates, including their validity and security analysis results.

### v2

@@ 7.4. Changes in Coding Behavior
7.5. Generalizability to C
The transferability of the the findings obtained for Python regarding the impact of different prompting techniques was tested on C, as mentioned in Section 5.4. The results from the CodeQL evaluations are presented in Table 6. As in the case of Python (Section 6.3), we selected a sample of results (49 cases, which account for 10% of the results) from CodeQL for manual inspection. After conducting a reliability agreement test, we obtained a Kappa value of 0.82 averaged over the results of the 7 prompting techniques from GPT-4, suggesting that the results for C code from CodeQL are highly reliable.
It can be seen that all the prompt templates that incorporated some form of security specifications showed improved results compared to the baseline prompt (which did not include any security-related instructions). Hence, this further highlights the importance of including explicit security specifications in prompts when generating code using LLMs. Among all the techniques, RCI produced C code with the lowest weakness count, rate, and density, consistent with the findings on Python. This, in principle, suggests its applicability across different programming languages. However, unlike the Python results, where the persona/memetic proxy technique consistently performed the worst, the CWE-specific template using the zero-shot technique performed the worst in C (of course, aside from the baseline prompt). Overall, the findings from the C language experiment suggest that while the RCI technique remains the most effective, the relative performance of other prompting techniques may vary depending on the programming language.
Further data from this experiment including the LLM-generated C code files, the CodeQL responses and the type of security weaknesses detected in the code are provided in the replication package.
Table 6. The results of validity and security analysis of C code generated by GPT-4 using the 7 prompt templates. The count is the total number of security weaknesses detected by CodeQL, rate is the average number of security weaknesses per code and density is the average number of security weaknesses per LOC.
GPT-4
Prompt Type
# valid code
# LOC Security Weaknesses
MIN MAX Avg. Count Rate Density
baseline (0-shot) 67 6 91 27.88 19 0.283 0.009
naive-secure (0-shot) 67 8 86 35.98 15 0.223 0.005
CWE-specific (0-shot) 67 24 95 41.5 19 0.283 0.006
comprehensive (0-shot) 67 17 95 50.97 13 0.194 0.004
zero-shot CoT 67 8 81 36.95 12 0.179 0.004
RCI 67 13 157 56.80 11 0.164 0.002
persona/memetic proxy 67 21 73 40.29 11 0.164 0.004
Actionable Takeaways: When using LLMs or LLM-powered tools like ChatGPT or Copilot, which enable natural language user interaction in a software development environment, the following considerations must be taken into account: (1) Using RCI is preferable over the other techniques studied in this work, as RCI can largely improve the security of the generated code (up to an order of magnitude w.r.t weakness density) even when applied with just 2 iterations. This technique has stayed valuable over several versions of the LLM models, and, hence, there is an expectation that it will stay valid in the future as well. (2) As of today, state-of-the-art LLM provides better results in terms of weakness density when used for C generation w.r.t Python, and therefore it should be used with more caution in the latter case. (3) Nevertheless, the use of RCI might bring the defect density in Python to the same ballpark as in C. Therefore, this might justify the use of a slightly more complex prompting technique for software development in practice. (4) In cases where multi-step techniques like RCI are not feasible, using simple zero-shot prompting with templates similar to comprehensive prompts, that specify well-established secure coding standards, can provide comparable results in relation to more complex techniques. (5) Even with the use of secure prompting techniques, coding tasks involving web application development and OS command execution can still result in security weaknesses due to their complex nature. Special attention should be given to such tasks, particularly for vulnerabilities related to CWE-259:Hard-coded Passwords (or more broadly CWE-798: Hard-coded Credentials), CWE-94 and CWE-78:OS Command Injection. (6) Static analysis tools like Bandit and CodeQL seem to be able to detect the issues that LLMs may overlook (such as CWE-259, CWE-94, CWE-78). Hence, it is still strongly advised to use these tools in tandem with LLMs in the development pipeline.
8. Impact of Data Leakage
Evaluating widely-used closed-source LLMs such as the ones used in this study poses the risk of data leakage (Ye et al., 2023). Data leakage (also referred to as data contamination) (Balloccu et al., 2024)occurs when models have prior exposure to the benchmark datasets used for their evaluation, which can lead to false estimations of their capabilities. Balloccu et al. (Balloccu et al., 2024) identified two forms of data leakage: direct and indirect.
Direct data leakage occurs when evaluation data is already included in a model’s training data. Considering that the OpenAI models used in this study are closed-source and their training data is undisclosed, this poses a potential concern. The LLMSecEval dataset used in this study, comprising NL code generation prompts paired with secure Python implementations, was formally published in May 2023. However, its corresponding GitHub public repository was created in January 2023. The OpenAI documentation [^5] https://platform.openai.com/docs/models states that GPT-3 (text-davinci-002) and GPT-3.5 (gpt-3.5-turbo) were trained on data up to September 2021, eliminating the risk of direct data leakage for these models. However, GPT-4 (gpt-4-1105-preview) was trained on data up to April 2023, suggesting a slight risk of leakage since the GitHub repository predates this one. To verify the extent of direct data leakage in the code generated by GPT-4, we performed a leakage test on the Python code generated with the RCI technique, the best-performing prompting technique in our experiments. For this, we employed the Dolos toolkit (Maertens et al., 2022), which is a source code plagiarism detection tool. Dolos works by tokenizing and canonicalizing programs into Abstract Syntax Tree (AST) representations to calculate a similarity score that captures semantic-level similarity through k-gram matching between the source and target programs. This tool has been used in several studies for quantifying contamination in code generated by LLMs (Riddell et al., 2024; Yu et al., 2023). The average similarity score obtained for all the code generated using RCI technique and the secure implementation present in the LLMSecEval dataset is 0.075. Yu et al. (Yu et al., 2023) in their study shows that a similarity score greater than 0.5 indicates potential plagiarism. Hence, this suggests that the impact of the direct data leakage/contamination in our results is very minimal.
On the other hand, indirect leakage happens when models learn from user interactions via Reinforcement Learning Through Human Feedback (RLHF). The OpenAI documentation [^6] https://help.openai.com/en/articles/5722486-how-your-data-is-used-to-improve-model-performance states that only data from web interface interactions, not from API usage, is utilized for this purpose. Since this study exclusively relied on API interactions, there is no risk of indirect data leakage.
@@ 9. Threats to Validity
Construct Validity. The validity analysis of code responses generated by all LLMs was conducted by a single author, potentially introducing biases in the evaluation process. Nonetheless, efforts were made to mitigate such biases by explicitly outlining the criteria for assessing the validity of code snippets, as detailed in Section 5.2. We also acknowledge that the prompting techniques underwent evaluation using prompt templates created by us. These generated templates might have influenced the results obtained for each technique from the LLMs. However, attention was given to crafting the templates, adhering closely to the design and examples that demonstrated optimal results in the respective papers that introduced these techniques.
External Validity. This study was conducted only using the OpenAI models. As previously stated in 5.1, this decision was made due to the popularity of these models in the prompt engineering literature and their demonstrated proficiency in handling coding tasks articulated in natural language, as identified during a preliminary model selection examination by us. It is also worth mentioning that, we focused on prompting techniques that do not rely on demonstrative examples. This choice stemmed from a user study (Perry et al., 2023) which highlighted that users predominantly interacted with AI assistants using natural language coding task specifications or instructions, without supplying demonstrative examples. We also note that the responses generated by the LLMs were not subjected to a randomness check. Since the code validity analysis was conducted manually, generating and evaluating multiple random responses from the LLMs was not feasible. However, we used 150 coding tasks to assess the impact of each prompting technique, which helps smooth out the fluctuation in results caused by randomness to some extent. Furthermore, the LLMSecEval dataset contains NL prompts for only 18 out of Top 25 CWEs of the year 2021. However, 15 out of the 18 CWEs considered in this study have retained their position on the Top 25 list of 2022 and 2023 proving the continued relevance and significance of our research findings. Additionally, while our agreement test between manual and Bandit security analyses demonstrates Bandit’s reliability to an extent, we acknowledge that it is not perfect. Some specific security weaknesses may have gone undetected. Nevertheless, Bandit proved sufficiently reliable for comparing the relative effectiveness of different prompting techniques.
@@ 10. Conclusion
Our analysis reaffirms the prevalence of security weaknesses in code generated by LLMs when prompted with NL instructions, with significant challenges stemming from CWE-78, CWE-259, CWE-94, and CWE-330. Among the prompting techniques investigated, RCI, a refinement-based approach, exhibited notable effectiveness in preventing security weaknesses in LLM-generated code. Particularly noteworthy was its performance with GPT-4, where it reduced the average weakness density by 77.5% compared to baseline prompting that includes no security specifications. Although RCI demonstrated the highest performance, to the extent of our knowledge, this technique has not been applied for secure code generation in existing literature. This highlights the need for additional research to investigate refinement-based methods, like RCI, which leverage self-critiquing and improvement capabilities of LLMs to enhance security in LLM-generated code. Zero-shot prompting yielded surprisingly favorable outcomes considering its straightforward nature, performing better than zero-shot CoT and persona/memetic proxy yet falling short of RCI. However, zero-shot prompting holds promise due to its simplicity and relative performance, provided an optimal prompt can be identified.
@@ 11. Replication Package
All data collected and generated in this study are available in https://figshare.com/s/195a75d8c4dd86816223. This repository contains the results of the literature review along with the information on the prompting techniques that were removed from our final selection and the rationale behind this exclusion. Furthermore, the repository contains the code generated by all 3 LLMs for the 7 prompt templates, including their validity and security analysis results.
@@
Appendix A Manual Verification of SAST Tool Results
Below, we detail the full process followed to manually verify the security analysis results of the SAST tools (Bandit and CodeQL).
A.1. Step 1: Manual Security Weakness Detection
We began by manually reviewing the code snippets (for the randomly selected tasks) independently of the SAST tool results to identify any of the 18 security weaknesses documented in LLMSecEval. This process involved consulting the official CWE pages on the MITRE website, specifically examining the provided potential mitigations and examples of vulnerable code for each weakness. A code snippet was flagged as vulnerable if it lacked the recommended mitigation techniques to prevent the vulnerability or closely resembled the examples of vulnerable code. Consider the example of CWE-78: OS Command Injection. One of the mitigation techniques outlined on the CWE page is as follows [^7] https://cwe.mitre.org/data/definitions/78.html:
Output encoding: While it is risky to use dynamically-generated query strings, code, or commands that mix control and data together, sometimes it may be unavoidable. Properly quote arguments and escape any special characters within those arguments. The most conservative approach is to escape or filter all characters that do not pass an extremely strict allowlist (such as everything that is not alphanumeric or white space). If some special characters are still needed, such as white space, wrap each argument in quotes after the escaping/filtering step.
Following this mitigation technique, we review the code for any instances of dynamically generated strings, code, or commands that lack validation and encoding. When such cases were identified, they were flagged as susceptible to CWE-78. The example snippet in Listing 1 shows a command that is executed with user-provided information without any validation or encoding, thus leading to CWE-78. Each code was checked for any of the 18 weaknesses following this approach manually.
1import subprocess
2import sys
3
4def get_user_info(username):
5
6 command = f"cat /etc/passwd | grep {username}"
7 result = subprocess.run(command, shell=True, capture_output=True, text=True)
8
9 if result.returncode == 0 and result.stdout:
[…]
11 else:
12 print(f"No user information found for: {username}")
13
14if __name__ == "__main__":
15 if len(sys.argv) != 2:
16 print("Usage: python script.py <username>")
17 else:
18 username = sys.argv[1]
19 get_user_info(username)
Listing 1: An example Python implementation to demonstrate CWE-78. The code takes a user input to construct a command to be executed.
A.2. Step 2: Verification of the Additional CWEs Reported by the SAST Tools
While running Bandit and CodeQL, we did not restrict the scope to the 18 weaknesses in LLMSecEval. This is to ensure that we also capture other weaknesses that may be present in the code snippets as our goal was to evaluate the security of the LLM-generated code as a whole. This has led to the detection of weaknesses that are not on our list such as CWE-94, CWE-259, CWE-327 and CWE-330. Few of these weaknesses had a parent/child relationship with the CWE IDs on our list. For instance, CWE-259888https://cwe.mitre.org/data/definitions/259.html is a child of CWE-798999https://cwe.mitre.org/data/definitions/798.html. These were identified as the same weakness in both manual and SAST tool analysis results. The remaining ones reported by the tool were manually inspected by checking the recommendations from MITRE for the corresponding weakness (as in Step 1). The verified weaknesses were added to manual inspection’s results.
A.3. Step 3: Agreement Test
The final outcomes of the manual security analysis from Step 2 and the SAST tool results were subjected to a reliability agreement test to evaluate the consistency between the two. This test was conducted using a weighted Cohen’s Kappa and the results were reported, accordingly.
Appendix B Security Analysis Using CodeQL
As mentioned in section 5.3 and 6.3, we also performed security evaluations of the LLM-generated Python code using CodeQL (in addition to Bandit). Table 7 shows the results of the CodeQL evaluation. The table follows the same structure as that of Table 3. Compared to Bandit, CodeQL identified fewer weaknesses in LLM-generated code. Despite the difference in the number of detected weaknesses, the results from CodeQL follow a similar pattern to those of Bandit.
Best Performer
Also according to CodeQL, the RCI prompting technique consistently produces the lowest weakness rate and density in code generated by GPT-3.5 and GPT-4, highlighting its superiority over other techniques.
Worst Performer
Aside from the baseline prompting template, the persona/memetic proxy technique results in the highest weakness rate in code generated by both GPT-3.5 and GPT-4, confirming the findings from Bandit.
The complete evaluation results generated by CodeQL are provided in the replication package.
Table 7. The results of validity and security analysis of Python code generated by the 3 LLMs using the 7 prompt templates. The count is the total number of security weaknesses detected by CodeQL, rate is the average number of security weaknesses per code and density is the average number of security weaknesses per LOC.
GPT-3
Prompt Type
# valid code
# LOC Security Weaknesses
MIN MAX Avg. Count Rate Density
baseline (0-shot) 131 2 80 11.175 53 0.404 0.024
naive-secure (0-shot) 123 2 31 10.691 19 0.153 0.013
CWE-specific (0-shot) 124 3 65 13.846 22 0.177 0.009
comprehensive (0-shot) 120 4 56 15.991 27 0.225 0.015
zero-shot CoT 126 3 32 10.753 32 0.253 0.021
RCI 125 2 84 20.960 28 0.224 0.012
persona/memetic proxy 137 5 76 15.875 31 0.226 0.014
GPT-3.5
Prompt Type
# valid code
# LOC Security Weaknesses
MIN MAX Avg. Count Rate Density
basic (0-shot) 145 3 38 13.889 67 0.462 0.028
naive-secure (0-shot) 147 3 55 16.374 49 0.333 0.020
CWE-specific (0-shot) 139 3 58 18.733 54 0.038 0.020
comprehensive (0-shot) 141 5 65 20.680 58 0.375 0.023
zero-shot CoT 140 3 42 14.357 47 0.335 0.018
RCI 138 5 65 23.543 35 0.253 0.008
persona/memetic proxy 141 2 42 12.970 57 0.404 0.042
GPT-4
Prompt Type
# valid code
# LOC Security Weaknesses
MIN MAX Avg. Count Rate Density
basic (0-shot) 144 3 39 16.990 91 0.631 0.032
naive-secure (0-shot) 149 5 65 21.738 77 0.516 0.020
CWE-specific (0-shot) 145 6 81 28.379 68 0.468 0.014
comprehensive (0-shot) 147 3 66 26.891 64 0.435 0.013
zero-shot CoT 146 3 68 22.246 51 0.349 0.016
RCI 143 3 94 39.902 48 0.335 0.006
persona/memetic proxy 147 3 50 19.319 77 0.523 0.024